how to prepare 5 L of 0.N hcl from 37% hydrochloric acid stock with density 1.19kg/L


1
Expert's answer
2021-02-12T04:59:49-0500

molar mass of HCl = 36.411

volume = 5L

HCl density = 1.19

grams of compound needed =

(0.1N)×36.411× 5 = 18.20g

Volume of concentrated acid needed =

18.20/(0.37×1.19) = 41.33 ml

So , 41.33 ml of 37% concentrated HCl is needed to prepare 0.1M HCl